OpenStudy (anonymous):

Find the point of inflection and the local maximum and minimum values of the function y=2sin(x)+sin^2(x) in the domain of (0, 2pi) This is what I have so far: y=2sin(x)+sin^2(x) y'=2cos(x)+2sin(x)cos(x) y'=2cos(x)+sin(2x) y''=-2sin(x)+2cos(2x) -2sin(x)+2cos(2x)=0 -sin(x)+cos(2x)=0 cos(2x)=sin(x) 1-sin^2(x)=sin(x) 2sin^2(x)+sin(x)-1=0 sin(x)=-1 & sin(x)=1/2 Please help!

OpenStudy (anonymous):

Find the zeros of the second derivative. On a trig function like this one, those will be the points of inflection of the function. The zero of the second derivative represents the places where the curvature is going from convex upward to convex downward, or visa-versa.

OpenStudy (anonymous):

Yes, so I got that I think, is it right?

OpenStudy (anonymous):

I follow it until the step where cos (2x) becomes 1-sin^2(x). I don't think this is justified.

OpenStudy (anonymous):

So what should I do?

OpenStudy (callisto):

cos2x = 1-2sin^2x ...

OpenStudy (anonymous):

Well, one solution is x= pi/6 in that case \[\cos(\frac{\pi}{3})=\sin(\frac{\pi}{6})\]

OpenStudy (anonymous):

Which I think is true.

OpenStudy (anonymous):

So if I corrected it to cos2x = 1-2sin^2x what would my final answer be?

OpenStudy (callisto):

In your workings : cos(2x)=sin(x) 1-sin^2(x)=sin(x) <- incorrect 2sin^2(x)+sin(x)-1=0 <- correct sin(x)=-1 & sin(x)=1/2 <- correct

OpenStudy (anonymous):

Oh, so its a typo

OpenStudy (anonymous):

I haven't taught trig in quite a while.... what is the identity that justifies this? I'm sorry to be intransigent, but I'm not seeing it.

OpenStudy (anonymous):

To find the inflection point you said we need to equate the equation to 0, ive done that and got sin(x)= -1, 1/2, are those the inflection points?

OpenStudy (callisto):

@AnimalAin it's cos2x = cos^2x - sin2x = 1-2sin^2x = 2cos^2x-1

OpenStudy (anonymous):

sin x = 1/2 corresponds to the answer I got earlier. I don't think arcsin -1 meets the specifications of the earlier condition (that would be cos 2x = sin x. Thanks Callisto; I always need to look up those double angle and half angle formulas; don't have my books handy here.

OpenStudy (anonymous):

The other postition of x that generates sin x = 1/2 is x= 5pi/6. Twice that position is in the fourth quadrant, and would give a negative value for cos 2x, so it couldn't meet the specification of the problem, either.

OpenStudy (callisto):

Not yet sin(x)= -1 or sinx = 1/2 x= 3pi /2 x = pi/6 or x=5pi /6 | x<pi/6 | pi/6 < x<5pi/6 | 5pi/6<x<3pi/2 | x>3pi/2 y''| + | - | + | + Therefore , inflection points are at x=pi/6, x=5pi/6 y=2sin(x)+sin^2(x) -(1) Put x = pi/6 into (1) y= 1.25 Put x=5pi/6 into (1) y= 1.25 So inflection points are (pi/6 , 1.25) and (5pi/6 , 1.25)

OpenStudy (callisto):

AnimalAin You're welcome

OpenStudy (anonymous):

How did you get the x= values?

OpenStudy (callisto):

sin 270degrees = -1 sin 30 degrees = 1/2 sin 150 degrees = 1/2 Use sin^(-1) 1/2 and sin^(-1) -1

OpenStudy (callisto):

x= sin^(-1) 1/2 or x = sin^(-1) -1

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

So how do we find the local max and mins?

OpenStudy (callisto):

Put y'=0 and solve x, then find the respective y value(s)

OpenStudy (anonymous):

I dont understand...

OpenStudy (callisto):

Which part?

OpenStudy (callisto):

@JoBo

OpenStudy (anonymous):

I'm sorry, I was confused at this end..... arcsin (-1)= 3pi/2 and cos(3pi)=-1

OpenStudy (anonymous):

put y'=0

OpenStudy (callisto):

@AnimalAin arcsin (-1)= 3pi/2 <- this is correct, but this is not the inflection point as the y'' is doesn't change, see the table i typed may help you understand, i hope. And I don't think I have mentioned ' cos(3pi)=-1' ... Sorry :(

OpenStudy (callisto):

@JoBo For a maximum/minimum point, the slope at that point =0 Therefore y'=0, that is 2cos(x)+sin(2x) =0 Solve x from here. Understand?

OpenStudy (anonymous):

Oh, thank you so much I finally understand! :) lol but I dont know how to solve for x...

OpenStudy (anonymous):

@callisto?

OpenStudy (callisto):

Sorry... was away for a while...

OpenStudy (anonymous):

Thats ok :)

OpenStudy (anonymous):

Can you teach me how to solve for x? for 2cos(x)+sin(2x) =0

OpenStudy (callisto):

2cos(x)+sin(2x) =0 2cos(x)+2sin(x)2cos(x) =0 2cosx (1+sin(x))=0 2cosx =0 or 1+sinx =0 cosx=0 or sinx =-1 x = pi/2 or x= 3pi/2 or x = 3pi/2 (repeated) x<pi/2 | pi/2 <x < 3pi/2 | x >3pi/2 y' | + | - | + (slope) / \ / So max is at x=pi/2 min is at 3pi/2 Work out the y values now

OpenStudy (callisto):

@JoBo

OpenStudy (anonymous):

Sorry about that

OpenStudy (anonymous):

Working out the y values will get me the coordinates for the max and min?

OpenStudy (callisto):

yes, put x value(s) into 2sin(x)+sin^2(x) to get the corresponding y-value(s)

OpenStudy (anonymous):

I still dont understand how you did this part... y=2sin(x)+sin^2(x) -(1) Put x = pi/6 into (1) y= 1.25 Put x=5pi/6 into (1) y= 1.25

OpenStudy (callisto):

You know I solve the x values for inflection point right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I dont understand how you got 1.25

OpenStudy (callisto):

you know the curve is a graph of y=2sin(x)+sin^2(x) , right?

OpenStudy (anonymous):

Yes

OpenStudy (callisto):

now you have x value, how would you get the y value?

OpenStudy (anonymous):

The x value being ie. pi/6 where do I plug that in?

OpenStudy (callisto):

when you see x, replace it by pi/6 in the function of the graph

OpenStudy (anonymous):

How would I input that into my calculator?

OpenStudy (callisto):

is it in radian or in degree mode?

OpenStudy (anonymous):

I can change it

OpenStudy (anonymous):

but yeah, degree mode

OpenStudy (callisto):

Sorry pi/6 = 30 degree Put 2sin(30)+sin^2(30) in your calculator, then it will give you the value 1.5

OpenStudy (anonymous):

I cant do sin^2(30) it does sin(30)^2

OpenStudy (anonymous):

@callisto?

OpenStudy (callisto):

Sorry.. I didn't notice your response.. sorry sin^2(30) = sin(30)^2

OpenStudy (anonymous):

When I do 2sin(30)+sin(30)^2 I get 1

OpenStudy (callisto):

Oopss.... it should be 1.25!! How silly me!!1

OpenStudy (anonymous):

Yea, but I get 1

OpenStudy (callisto):

try again~~~ 2sin(30)+sin(30)^2 = 2(1/2) + (1/2)^2 = 1+0.25 = 1.25

OpenStudy (anonymous):

Oh yea youre right

OpenStudy (callisto):

Got it?!

OpenStudy (anonymous):

For this post: 2cos(x)+sin(2x) =0 2cos(x)+2sin(x)2cos(x) =0 2cosx (1+sin(x))=0 2cosx =0 or 1+sinx =0 cosx=0 or sinx =-1 x = pi/2 or x= 3pi/2 or x = 3pi/2 (repeated) x<pi/2 | pi/2 <x < 3pi/2 | x >3pi/2 y' | + | - | + (slope) / \ / So max is at x=pi/2 min is at 3pi/2 Work out the y values now I put the x values back into which equation?

OpenStudy (callisto):

Original one. Whenever you need to find the corresponding y-value of x in the graph, put x into the original one , that is not y' nor y''

OpenStudy (anonymous):

What does finding the y values do?

OpenStudy (anonymous):

So I do 2cos(90)+sin(180)?

OpenStudy (anonymous):

@Callisto?

OpenStudy (callisto):

nope :S then x value input should be the same every time :S

OpenStudy (anonymous):

Im confused...

OpenStudy (anonymous):

are you talking about sin(2x)? I just multiplied it

OpenStudy (callisto):

I see... but the original equation is in sin only... no cos as i see...

OpenStudy (anonymous):

I did 2cos(x)+sin(2x) =0

OpenStudy (anonymous):

Are you talking about 2sin(x)+sin^2(x) ?

OpenStudy (callisto):

yes.... this is the original one :S

OpenStudy (anonymous):

I get 0 for the first one and -1 for the second one

OpenStudy (anonymous):

What do I do with these numbers?

OpenStudy (callisto):

this is the corresponding values of x the coord = (pi/2, 0) for max (3pi/2 , -1) for min Assume you've not made any mistakes

OpenStudy (anonymous):

I see, THANK YOU SO MUCH Callisto!!! <333

OpenStudy (callisto):

welcome!!!